# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 1 Introduction
From a more physical point of view, a treatment non-invertible topological lines in 2D theories as symmetries was pioneered in [39, 40, 41] and further discussed in [42, 43]. The relationship between the 3D bulk TQFT and 2D boundary CFT is a foundational example of the general paradigm of a bulk topological field theory controlling the generalized symmetry of the boundary theory [18, 44, 45, 46, 47, 48]. The general idea of gauging and condensing non-invertible symmetries has been explored in [49, 50, 51, 52, 53, 54] and the relationship between bulk and boundary non-invertible gauging has been utilized in [55, 56, 57, 58, 59, 60]. Finally, recent discussions of gauging non-invertible symmetries in 2D CFTs, closely related to our analysis below, include in particular [61, 62]. Below we often make use of the language of generalized symmetries, interchangeably using non-abelian and non-invertible, as well as gauge and condense.
@@ 1.1 An Invitational Example with Fibonacci Anyons
We now use that $(G_{2})_{1}\times(G_{2})_{-1}$ is a Drinfeld center, i.e., it is of the form $G_{k}\times G_{-k}$. In particular for any such theory it is known that one can gauge/condense all the anyons and obtain a trivial theory [36, 33, 67]. The novelty here is that such condensation is necessarily non-abelian. Concretely then, we write
@@ 1.2 Summary of Selected Results
which we explicitly check for $k=2$ in Section 4.2 (for $k=1$ this expression may be checked by the three-step gauging rule).
@@ 2 Cosets, Interfaces, and Bulk-Boundary Correspondence
A subtlety in the construction above is the generic appearance of multiple copies of the vacuum and of many copies of the same chiral or Virasoro primary in the partition function. Relatedly, not all branching functions are non-zero, so the naive modular covariance of the branching functions in general requires further analysis.111 Historically, these confusions led to the search for methods to remove such a degeneracy, such as the so-called “identification current method” or “fixed point resolutions” that made the final result into a CFT with a single vacuum, a non-degenerate modular S-matrix, etc. The result of such a procedure is what in some older literature is known as “the” coset CFT (See for instance [15, 71, 72, 73]). In modern terms, we recognize the vacuum degeneracy as the presence of a topological sector coupled to gapless, CFT degrees of freedom. In some circumstances it is desirable to remove this degeneracy by, roughly speaking, “gauging away” this topological sector resulting in a CFT with a unique vacuum. In general, we should therefore differentiate between two possible notions of cosets:
@@
Figure 5: In a topological boundary condition not all lines of the bulk theory end perpendicularly to the boundary. Only those generating a Lagrangian algebra can end, resulting in a topological local operator at the junction. When $Z$ is abelian as it is assumed in this section, the Lagrangian algebra is moreover a Lagrangian subgroup. Under parallel fusion, those lines in the Lagrangian subgroup become invisible at the boundary.
@@ 3 Dualities via Non-Invertible Anyon Condensation
In the context of the GKO coset construction, it is also known that there are exceptions to the methods mentioned in footnote 1 to construct CFTs with single vacuum from the “naive” partition function (2.3). In these exceptions there are additional selection rules (i.e.  vanishing branching functions) and field identifications that are not a consequence of group-theoretical selection rules and thus cannot be submitted to the “field identification method” [72] mentioned above to find a CFT with single vacuum and non-degenerate $S$-matrix. Importantly, again the conformal embeddings appear as part of such exceptions (see [15, 91]). We stress that if the conformal embedding has two factors in the denominator we mean the full quotient appears as an exception, and not a standard coset based on just one of the factors. Furthermore, it is known that a handful of gapless cosets also share this feature. In the literature, these cosets have been referred to as “Maverick cosets,” and the known examples have been constructed long ago in [91, 20, 21, 71]. In some sense, Maverick cosets are then a hybrid in between the standard cosets and the conformal embeddings. To the extent of the authors’ knowledge, there is no known classification of Maverick cosets.
@@
Then, by this duality, in the bulk theory $(G_{2})_{1}\times SU(2)_{-3}$ we are allowed to take as a boundary condition a canonical $SU(2)_{1}$ RCFT boundary condition for the $SU(2)_{1}$ factor, but the purely topological boundary condition for the factor $(G_{2})_{1}\times(G_{2})_{-1}$ given by the diagonal Lagrangian algebra. Recall this type of topological boundary is given by the $G_{k}/G_{k}$ 2D TQFT, which has as many topological local operators as the $G_{k}$ WZW theory has chiral primaries, with the same fusion rules as the $G_{k}$ WZW theory. The possibility of choosing this “combined CFT-Topological” boundary condition explains both the degeneracy of two in the torus partition function (3.6) as well as the precise Fibonacci TQFT factor in (3.7), since $(G_{2})_{1}$ has Fibonacci fusion rules.
@@ 4 Explicit Checks via Non-Abelian Anyon Condensation
First, we summarize some consistency conditions that will allow us to pin down the resulting theory after non-abelian anyon condensation. We follow the heuristic rules of [23] for this purpose. Non-abelian anyon condensation was formalized in [38], but the rules of [23] are more useful for practical calculations. Following the logic of the three-step gauging procedure reviewed in Section 1, we aim to obtain the spectrum and topological spins of the anyons in the gauged theory without working out the full MTC data. The rules of [23] enable this approach. In successive subsections, we consider many explicit examples of interest. In the following, we refer as parent theory to the original theory before condensation/gauging, and as child theory to the one obtained after condensation/gauging. For obvious reasons, we also refer to the parent and child as uncondensed and condensed theories, respectively. As in the case of standard gaugings of abelian anyons (one-form global symmetries) only bosons can condense, reflecting the fact that the topological spins capture the ‘t Hooft anomalies [18, 11, 19]. [^10] Relatedly, commutative algebras in a MTC decompose necessarily in terms of simple anyons with trivial topological spin. In particular then, special symmetric commutative Frobenius algebras Frobenius algebras decompose in terms of simple anyons with trivial topological spin.
@@
Line label
Quantum Dimension
Conformal Weight
0
$d_{0}=1$
$h_{0}=0$
$\phi$
$d_{\phi}=\frac{1+\sqrt{5}}{2}$
$h_{\phi}=2/5$
@@
Line label
Quantum Dimension
Conformal Weight
0
$d_{0}=1$
$h_{0}=0$
1
$d_{1}=\frac{1+\sqrt{5}}{2}$
$h_{1}=3/20$
2
$d_{2}=\frac{1+\sqrt{5}}{2}$
$h_{2}=2/5$
3
$d_{3}=1$
$h_{3}=3/4$
@@ 4.2 Checking the $SU(2)_{1}\times SU(2)_{2}\cong\big{(}M(5,4)\times SU(2)_{3}\big{)}/\mathcal{A}$ Duality
Line label
Quantum Dimension
Conformal Weight
0
$d_{0}=1$
$h_{0}=0$
$\varepsilon$
$d_{\varepsilon}=\frac{1+\sqrt{5}}{2}$
$h_{\varepsilon}=1/10$
$\varepsilon^{\prime}$
$d_{\varepsilon^{\prime}}=\frac{1+\sqrt{5}}{2}$
$h_{\varepsilon^{\prime}}=3/5$
$\varepsilon^{\prime\prime}$
$d_{\varepsilon^{\prime\prime}}=1$
$h_{\varepsilon^{\prime\prime}}=3/2$
$\sigma$
$d_{\sigma}=\sqrt{2}\big{(}\frac{1+\sqrt{5}}{2}\big{)}$
$h_{\sigma}=3/80$
$\sigma^{\prime}$
$d_{\sigma^{\prime}}=\sqrt{2}$
$h_{\sigma^{\prime}}=7/16$
@@
Line label
Quantum Dimension
Conformal Weight
(0,0)
$d_{(0,0)}=1$
$h_{(0,0)}=0$
(0,1)
$d_{(0,1)}=1$
$h_{(0,1)}=1/4$
(1,0)
$d_{(1,0)}=\sqrt{2}$
$h_{(1,0)}=3/16$
(1,1)
$d_{(1,1)}=\sqrt{2}$
$h_{(1,1)}=7/16$
(2,0)
$d_{(2,0)}=1$
$h_{(2,0)}=1/2$
(2,1)
$d_{(2,1)}=1$
$h_{(2,1)}=3/4$
@@ 4.3 Checking the $U(1)^{\mathrm{Orb}}_{6}\cong(SU(4)_{2}\times Spin(4)_{-4})/\mathcal{A}$ Duality
Line label
Quantum Dimension
Conformal Weight
$\mathbf{1}$
$d_{\mathbf{1}}=1$
$h_{\mathbf{1}}=0$
$\overline{\mathbf{10}}$
$d_{\overline{\mathbf{10}}}=1$
$h_{\overline{\mathbf{10}}}=3/4$
$\mathbf{20}^{\prime}$
$d_{\mathbf{20}^{\prime}}=1$
$h_{\mathbf{20}^{\prime}}=1$
$\mathbf{10}$
$d_{\mathbf{10}}=1$
$h_{\mathbf{10}}=3/4$
$\overline{\mathbf{4}}$
$d_{\overline{\mathbf{4}}}=\sqrt{3}$
$h_{\overline{\mathbf{4}}}=5/16$
$\mathbf{20}$
$d_{\mathbf{20}}=\sqrt{3}$
$h_{\mathbf{20}}=13/16$
$\overline{\mathbf{20}}$
$d_{\overline{\mathbf{20}}}=\sqrt{3}$
$h_{\overline{\mathbf{20}}}=13/16$
$\mathbf{4}$
$d_{\mathbf{4}}=\sqrt{3}$
$h_{\mathbf{4}}=5/16$
$\mathbf{15}$
$d_{\mathbf{15}}=2$
$h_{\mathbf{15}}=2/3$
$\mathbf{6}$
$d_{\mathbf{6}}=2$
$h_{\mathbf{6}}=5/12$
@@
Line label
Quantum Dimension
Conformal Weight
0
$d_{0}=1$
$h_{0}=0$
1
$d_{1}=\sqrt{3}$
$h_{1}=1/8$
2
$d_{2}=2$
$h_{2}=1/3$
3
$d_{3}=\sqrt{3}$
$h_{3}=5/8$
4
$d_{4}=1$
$h_{4}=1$
@@
Line label
Quantum Dimension
Conformal Weight
0
$d_{0}=1$
$h_{0}=0$
1
$d_{1}=1$
$h_{1}=3/4$
2
$d_{2}=1$
$h_{2}=1$
3
$d_{3}=1$
$h_{3}=3/4$
4
$d_{4}=2$
$h_{4}=1/12$
5
$d_{5}=2$
$h_{5}=1/3$
6
$d_{6}=\sqrt{3}$
$h_{6}=1/16$
7
$d_{7}=\sqrt{3}$
$h_{7}=9/16$
8
$d_{8}=\sqrt{3}$
$h_{8}=9/16$
9
$d_{9}=\sqrt{3}$
$h_{9}=1/16$
@@
Line label
Quantum Dimension
Conformal Weight
(0,0)
$d_{(0,0)}=1$
$h_{(0,0)}=0$
(0,4) $\cong$ (4,0)
$d_{(0,4)}=1$
$h_{(0,4)}=1$
(2,0) $\cong$ (2,4)
$d_{(2,0)}=2$
$h_{(2,0)}=1/3$
(1,1) $\cong$ (3,3)
$d_{(1,1)}=3$
$h_{(1,1)}=1/4$
(1,3) $\cong$ (3,1)
$d_{(1,3)}=3$
$h_{(1,3)}=3/4$
(0,2) $\cong$ (4,2)
$d_{(0,2)}=2$
$h_{(0,2))}=1/3$
$(2,2)_{1}$
$d_{(2,2)_{1}}=2$
$h_{(2,2)_{1}}=2/3$
$(2,2)_{2}$
$d_{(2,2)_{2}}=2$
$h_{(2,2)_{2}}=2/3$
@@ 5.1 Revisiting Classical Conformal Embeddings
where $\mathcal{A}_{N,k}$ is some Frobenius algebra object (for brevity, on the rest of this section when we write a letter in calligraphic font we always mean some appropriate Frobenius algebra object that we can use to gauge, and from now on we will not remark on what such objects stand for). The previous statement has been discussed in many mathematical references [22, 67, 98, 99, 38, 36], but in our context it is most quickly understood from the coset inversion formula (3.17) with $\mathcal{M}=SU(Nk)_{1}$, $\mathcal{C}=SU(k)_{N}$ and $\mathcal{M}^{\prime}=SU(N)_{k}$. In this sense, (5.2) is nothing but the standard form of cosets, Eqn. (3.18) of the coset inversion formulas, while (5.3) is “the parent statement” Eqn. (3.17) translated to a form more akin to physics, and for the particular example at hand. Clearly, the unitary groups play no role in the previous argument, and as such the same would hold for any other conformal embeddings. We explore this in the next sections below.

### v2

@@ 1 Introduction
From a more physical point of view, a treatment of non-invertible topological lines in 2D theories as symmetries was pioneered in [39, 40, 41] and further discussed in [42, 43]. The relationship between the 3D bulk TQFT and 2D boundary CFT is a foundational example of the general paradigm of a bulk topological field theory controlling the generalized symmetry of the boundary theory [18, 44, 45, 46, 47, 48]. The general idea of gauging and condensing non-invertible symmetries has been explored in [49, 50, 51, 52, 53, 54] and the relationship between bulk and boundary non-invertible gauging has been utilized in [55, 56, 57, 58, 59, 60]. Finally, recent discussions of gauging non-invertible symmetries in 2D CFTs, closely related to our analysis below, include in particular [61, 62]. Below we often make use of the language of generalized symmetries, interchangeably using non-abelian and non-invertible, as well as gauge and condense.
@@ 1.1 An Invitational Example with Fibonacci Anyons
We now use that $(G_{2})_{1}\times(G_{2})_{-1}$ is a Drinfeld double, i.e., it is of the form $G_{k}\times G_{-k}$. In particular for any such theory it is known that one can gauge/condense all the anyons and obtain a trivial theory [36, 33, 67]. The novelty here is that such condensation is necessarily non-abelian. Concretely then, we write
@@ 1.2 Summary of Selected Results
which we explicitly check for $k=2$ in Section 4.2 (for $k=1$ this expression may be checked by the three-step gauging rule). Finally, more examples of duality involving non-abelian anyon condensation may be found from the classification of RCFTs with a small number of primaries and a small central charge studied in [69, 70] and associated cosets of holomorphic CFTs at $c=24$ [71]. [^1] We thank B. Rayhaun for bringing this to our attention.
@@ 2 Cosets, Interfaces, and Bulk-Boundary Correspondence
A subtlety in the construction above is the generic appearance of multiple copies of the vacuum and of many copies of the same chiral or Virasoro primary in the partition function. Relatedly, not all branching functions are non-zero, so the naive modular covariance of the branching functions in general requires further analysis.222 Historically, these confusions led to the search for methods to remove such a degeneracy, such as the so-called “identification current method” or “fixed point resolutions” that made the final result into a CFT with a single vacuum, a non-degenerate modular S-matrix, etc. The result of such a procedure is what in some older literature is known as “the” coset CFT (See for instance [15, 74, 75, 76]). In modern terms, we recognize the vacuum degeneracy as the presence of a topological sector coupled to gapless, CFT degrees of freedom. In some circumstances it is desirable to remove this degeneracy by, roughly speaking, “gauging away” this topological sector resulting in a CFT with a unique vacuum. In general, we should therefore differentiate between two possible notions of cosets:
@@
Figure 5: In a topological boundary condition not all lines of the bulk theory end perpendicularly to the boundary. Only those generating a Lagrangian algebra can end (see Appendix B.1 for a precise mathematical definition), resulting in a topological local operator at the junction. When $Z$ is abelian as it is assumed in this section, the Lagrangian algebra is moreover a Lagrangian subgroup. Under parallel fusion, those lines in the Lagrangian subgroup become invisible at the boundary.
@@ 3 Dualities via Non-Invertible Anyon Condensation
In the context of the GKO coset construction, it is also known that there are exceptions to the methods mentioned in footnote 2 to construct CFTs with single vacuum from the “naive” partition function (2.3). In these exceptions there are additional selection rules (i.e.  vanishing branching functions) and field identifications that are not a consequence of group-theoretical selection rules and thus cannot be submitted to the “field identification method” [75] mentioned above to find a CFT with single vacuum and non-degenerate $S$-matrix. Importantly, again the conformal embeddings appear as part of such exceptions (see [15, 94]). We stress that if the conformal embedding has two factors in the denominator we mean the full quotient appears as an exception, and not a standard coset based on just one of the factors. Furthermore, it is known that a handful of gapless cosets also share this feature. In the literature, these cosets have been referred to as “Maverick cosets,” and the known examples have been constructed long ago in [94, 20, 21, 74]. In some sense, Maverick cosets are then a hybrid in between the standard cosets and the conformal embeddings. To the extent of the authors’ knowledge, there is no known classification of Maverick cosets.
@@
Then, by this duality, in the bulk theory $(G_{2})_{1}\times SU(2)_{-3}$ we are allowed to take as a boundary condition a canonical $SU(2)_{1}$ RCFT boundary condition for the $SU(2)_{1}$ factor, but the purely topological boundary condition for the factor $(G_{2})_{1}\times(G_{2})_{-1}$ given by the diagonal Lagrangian algebra (see Appendix B.1 for a precise mathematical definition). Recall this type of topological boundary is given by the $G_{k}/G_{k}$ 2D TQFT, which has as many topological local operators as the $G_{k}$ WZW theory has chiral primaries, with the same fusion rules as the $G_{k}$ WZW theory. The possibility of choosing this “combined CFT-Topological” boundary condition explains both the degeneracy of two in the torus partition function (3.6) as well as the precise Fibonacci TQFT factor in (3.7), since $(G_{2})_{1}$ has Fibonacci fusion rules.
@@ 4 Explicit Checks via Non-Abelian Anyon Condensation
First, we summarize some consistency conditions that will allow us to pin down the resulting theory after non-abelian anyon condensation. We follow the heuristic rules of [23] for this purpose. Non-abelian anyon condensation was formalized in [38], but the rules of [23] are more useful for practical calculations. Following the logic of the three-step gauging procedure reviewed in Section 1, we aim to obtain the spectrum and topological spins of the anyons in the gauged theory without working out the full MTC data. The rules of [23] enable this approach. In successive subsections, we consider many explicit examples of interest. In the following, we refer as parent theory to the original theory before condensation/gauging, and as child theory to the one obtained after condensation/gauging. For obvious reasons, we also refer to the parent and child as uncondensed and condensed theories, respectively. As in the case of standard gaugings of abelian anyons (one-form global symmetries) only bosons can condense, paralleling the fact that for abelian anyons the topological spins capture the ‘t Hooft anomalies [18, 11, 19]. [^11] Relatedly, commutative algebras in a MTC decompose necessarily in terms of simple anyons with trivial topological spin. In particular then, special symmetric commutative Frobenius algebras decompose in terms of simple anyons with trivial topological spin.
@@
Line label Quantum Dimension Conformal Weight
0 $d_{0}=1$ $h_{0}=0$
$\phi$ $d_{\phi}=\frac{1+\sqrt{5}}{2}$ $h_{\phi}=2/5$
@@ 0 $d_{0}=1$ $h_{0}=0$
Line label Quantum Dimension Conformal Weight
0 $d_{0}=1$ $h_{0}=0$
1 $d_{1}=\frac{1+\sqrt{5}}{2}$ $h_{1}=3/20$
2 $d_{2}=\frac{1+\sqrt{5}}{2}$ $h_{2}=2/5$
3 $d_{3}=1$ $h_{3}=3/4$
@@ 4.2 Checking the $SU(2)_{1}\times SU(2)_{2}\cong\big{(}M(5,4)\times SU(2)_{3}\big{)}/\mathcal{A}$ Duality
Line label Quantum Dimension Conformal Weight
0 $d_{0}=1$ $h_{0}=0$
$\varepsilon$ $d_{\varepsilon}=\frac{1+\sqrt{5}}{2}$ $h_{\varepsilon}=1/10$
$\varepsilon^{\prime}$ $d_{\varepsilon^{\prime}}=\frac{1+\sqrt{5}}{2}$ $h_{\varepsilon^{\prime}}=3/5$
$\varepsilon^{\prime\prime}$ $d_{\varepsilon^{\prime\prime}}=1$ $h_{\varepsilon^{\prime\prime}}=3/2$
$\sigma$ $d_{\sigma}=\sqrt{2}\big{(}\frac{1+\sqrt{5}}{2}\big{)}$ $h_{\sigma}=3/80$
$\sigma^{\prime}$ $d_{\sigma^{\prime}}=\sqrt{2}$ $h_{\sigma^{\prime}}=7/16$
@@ 0 $d_{0}=1$ $h_{0}=0$
Line label Quantum Dimension Conformal Weight
(0,0) $d_{(0,0)}=1$ $h_{(0,0)}=0$
(0,1) $d_{(0,1)}=1$ $h_{(0,1)}=1/4$
(1,0) $d_{(1,0)}=\sqrt{2}$ $h_{(1,0)}=3/16$
(1,1) $d_{(1,1)}=\sqrt{2}$ $h_{(1,1)}=7/16$
(2,0) $d_{(2,0)}=1$ $h_{(2,0)}=1/2$
(2,1) $d_{(2,1)}=1$ $h_{(2,1)}=3/4$
@@ 4.3 Checking the $U(1)^{\mathrm{Orb}}_{6}\cong(SU(4)_{2}\times Spin(4)_{-4})/\mathcal{A}$ Duality
Line label Quantum Dimension Conformal Weight
$\mathbf{1}$ $d_{\mathbf{1}}=1$ $h_{\mathbf{1}}=0$
$\overline{\mathbf{10}}$ $d_{\overline{\mathbf{10}}}=1$ $h_{\overline{\mathbf{10}}}=3/4$
$\mathbf{20}^{\prime}$ $d_{\mathbf{20}^{\prime}}=1$ $h_{\mathbf{20}^{\prime}}=1$
$\mathbf{10}$ $d_{\mathbf{10}}=1$ $h_{\mathbf{10}}=3/4$
$\overline{\mathbf{4}}$ $d_{\overline{\mathbf{4}}}=\sqrt{3}$ $h_{\overline{\mathbf{4}}}=5/16$
$\mathbf{20}$ $d_{\mathbf{20}}=\sqrt{3}$ $h_{\mathbf{20}}=13/16$
$\overline{\mathbf{20}}$ $d_{\overline{\mathbf{20}}}=\sqrt{3}$ $h_{\overline{\mathbf{20}}}=13/16$
$\mathbf{4}$ $d_{\mathbf{4}}=\sqrt{3}$ $h_{\mathbf{4}}=5/16$
$\mathbf{15}$ $d_{\mathbf{15}}=2$ $h_{\mathbf{15}}=2/3$
$\mathbf{6}$ $d_{\mathbf{6}}=2$ $h_{\mathbf{6}}=5/12$
@@
Line label Quantum Dimension Conformal Weight
0 $d_{0}=1$ $h_{0}=0$
1 $d_{1}=\sqrt{3}$ $h_{1}=1/8$
2 $d_{2}=2$ $h_{2}=1/3$
3 $d_{3}=\sqrt{3}$ $h_{3}=5/8$
4 $d_{4}=1$ $h_{4}=1$
@@
Line label Quantum Dimension Conformal Weight
0 $d_{0}=1$ $h_{0}=0$
1 $d_{1}=1$ $h_{1}=3/4$
2 $d_{2}=1$ $h_{2}=1$
3 $d_{3}=1$ $h_{3}=3/4$
4 $d_{4}=2$ $h_{4}=1/12$
5 $d_{5}=2$ $h_{5}=1/3$
6 $d_{6}=\sqrt{3}$ $h_{6}=1/16$
7 $d_{7}=\sqrt{3}$ $h_{7}=9/16$
8 $d_{8}=\sqrt{3}$ $h_{8}=9/16$
9 $d_{9}=\sqrt{3}$ $h_{9}=1/16$
@@
Line label Quantum Dimension Conformal Weight
(0,0) $d_{(0,0)}=1$ $h_{(0,0)}=0$
(0,4) $\cong$ (4,0) $d_{(0,4)}=1$ $h_{(0,4)}=1$
(2,0) $\cong$ (2,4) $d_{(2,0)}=2$ $h_{(2,0)}=1/3$
(1,1) $\cong$ (3,3) $d_{(1,1)}=3$ $h_{(1,1)}=1/4$
(1,3) $\cong$ (3,1) $d_{(1,3)}=3$ $h_{(1,3)}=3/4$
(0,2) $\cong$ (4,2) $d_{(0,2)}=2$ $h_{(0,2))}=1/3$
$(2,2)_{1}$ $d_{(2,2)_{1}}=2$ $h_{(2,2)_{1}}=2/3$
$(2,2)_{2}$ $d_{(2,2)_{2}}=2$ $h_{(2,2)_{2}}=2/3$
@@ 5.1 Revisiting Classical Conformal Embeddings
where $\mathcal{A}_{N,k}$ is some gaugable algebra object (for brevity, on the rest of this section when we write a letter in calligraphic font we always mean some appropriate algebra object that we can use to gauge, and from now on we will not remark on what such objects stand for). The previous statement has been discussed in many mathematical references [22, 67, 101, 102, 38, 36], but in our context it is most quickly understood from the coset inversion formula (3.17) with $\mathcal{M}=SU(Nk)_{1}$, $\mathcal{C}=SU(k)_{N}$ and $\mathcal{M}^{\prime}=SU(N)_{k}$. In this sense, (5.2) is nothing but the standard form of cosets, Eqn. (3.18) of the coset inversion formulas, while (5.3) is “the parent statement” Eqn. (3.17) translated to a form more akin to physics, and for the particular example at hand. Clearly, the unitary groups play no role in the previous argument, and as such the same would hold for any other conformal embeddings. We explore this in the next sections below.